# Multitask methods for predicting molecular properties from heterogeneous data

### v1

@@ II.2 $\Delta$-Learning
$\Delta$-learning performs Gaussian process regression to predict differences, $\Delta_{i}$, between observations obtained with two different methods for a given $\bm{X}_{i}$. If we retain our assumption that noise follows an independent, identically distributed Gaussian distribution, we have the model
$\bm{Y}_{ij_{1}}-\bm{Y}_{ij_{2}}$$\equiv$ $\ \Delta_{i}\ =\ f\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{i},$ (11)
$f$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}_{\Delta},\ k_{\Delta}(\bm{X},\bm{X}^{\prime})\big{)}.$
@@
where $j_{1}$ and $j_{2}$ indicate two prediction methods. An application of Eq. (10) yields a distribution on the regression function for $\Delta$. Suppose our ultimate interest is to predict $\bm{Y}$ at $\bm{X}_{*}$. Within the $\Delta$ framework, we can supply an observation from one method, say $\bm{Y}_{*j_{2}}$, in order to make a prediction for the value from the other method:
$\bm{Y}_{*j_{1}}\ \approx\ \Delta_{i}\ +\ \bm{Y}_{*j_{2}}.$
@@
Thus, $\Delta$-learning is suited to settings where calculations by method $j_{1}$ are both more accurate and more time consuming than calculations that use $j_{2}$.
@@
Consider a case where we want to build a model from an existing data set which includes predictions from CCSD(T) and several DFT methods. If the predictions all correspond to the same set of molecules, $\Delta$-learning models have the ability to leverage more of the data set than conventional GP regression models. Unfortunately, $\Delta$-learning is not-applicable in cases where predictions by different methods correspond to different molecular configurations. To use a DFT prediction for a molecule $\bm{X}_{i}$ to train a $\Delta$ model between CCSD(T) and DFT, we also require a CCSD(T) prediction for the same molecule. In the following subsections, we consider alternative approaches which model disparities in data sources and have the ability to use all available observations.
II.3 Multifidelity Method
@@
As a first step toward considering statistical models which train on heterogeneous data, we describe the multifidelity data fusion model. Kennedy and O’Hagan (2000) The multifidelity approach defines two regression functions: $f_{p}$,representing the predictions from some high fidelity model, and $f_{s}$, representing a low fidelity model. Kennedy and O’Hagan (2000) We assume the relationships
$\bm{Y}_{ip}$ $=$ $\ f_{p}\big{(}\bm{X}_{i}\big{)}\ +\varepsilon_{ip}$ (12)
$=$ $\ \rho f_{s}\big{(}\bm{X}_{i}\big{)}\ +\ \delta_{ps}\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{ip},$
$\bm{Y}_{is}$ $=$ $\ f_{s}\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{is}.$ (13)
@@
The term $\delta_{ps}$ captures the difference between the high fidelity model and the low fidelity model, scaled by hyperparameter, $\rho$. This disparity is endowed with a Gaussian process prior, as is the low fidelity regression function, $f_{s}$. The correlation, $\rho$, will be optimized before inference. As before, we assume that the additive noise terms, $\varepsilon_{ip}$ and $\varepsilon_{is}$, are independent and drawn from identical, centered Gaussian distributions. Now, suppose we want to predict our targets, $\bm{f}_{*}$, at the high fidelity level. We can determine that the joint distribution of our high and low fidelity observations and our targets has covariance
@@
where $\sigma^{2}$ is the variance of the noise distribution. The superscript of each covariance matrix, K, indicates which Gaussian process prior corresponds to the kernel function used in its construction, and the subscripts indicate the pairs of features that are compared. Bayesian inference yields an analytical posterior distribution, analogous to Eq. (10).
II.4 Multitask Method
“Multitasking” refers to a category of methods which consider several Gaussian process regression tasks and assume some relationship between these tasks. As with $\Delta$-learning or multifidelity fusion, multitasking can allow us to incorporate a larger dataset into our inference problem without just pooling the data into one observation vector. For instance, we can define a regression problem trained on CCSD(T) data as well as one trained on each DFA we consider. By modeling correlation between regression functions, we can use the DFA regression tasks to support prediction in the CCSD(T) task.
We can apply multitask regression to problems with $m$ tasks where, for each task $j\in\{1,\ldots,m\}$, we have observational data $\bm{Y}_{j}\in\mathbb{R}^{n}$. In practice, the method can be generalized, so that observation sets of different sizes are used for different tasks. Symmetric variants of multitask modeling treat all tasks equally. For instance, we may assume that all the data are modeled by a single Gaussian process whose covariance kernel is the Kronecker product of one kernel which relates the tasks and another which models relationships between data from the same task.Bonilla, Chai, and Williams (2008)
In many cases, however, the goal is to perform regression for a primary task, and all other tasks may be considered secondary tasks which provide data that may be useful in learning the primary task. For our own setting, CCSD(T) is a natural choice of data for the primary regression task because we would like to make predictions which match this method in accuracy. Each set of DFT predictions made using a different DFA could inform a secondary regression task which supports the primary task. Thus, we consider an asymmetrical model Leen, Peltonen, and Kaski (2012) in which secondary tasks are related through the primary task.
Suppose $\bm{Y}_{p}\in\mathbb{R}^{n_{p}}$ is the observation data for the primary task and $\bm{Y}_{s_{j}}\in\mathbb{R}^{n_{j}}$ is the data for the $s_{j}^{th}$ secondary task. As before, we assume that each task has its own regression function
$\bm{Y}_{ip}$ $=$ $\ f_{p}(\bm{X}_{i})\ +\ \varepsilon_{ip},$ (18)
$\bm{Y}_{is_{j}}$ $=$ $\ f_{s_{j}}(\bm{X}_{i})\ +\ \varepsilon_{is_{j}}\qquad\forall j=1,\dots,m.$ (19)
@@ II.4 Multitask Method
$f_{s_{j}}(\bm{X}_{i})$ $=\ \rho_{s_{j}}f_{p}(\bm{X}_{i})\ +\ \delta_{s_{j}}(\bm{X}_{i})\qquad\forall j=1,\dots,m.$ (20)
The specific component, $\delta_{s_{j}}$, of the secondary function serves to “explain away” behavior that is not captured by the shared component, $f_{p}$een, Peltonen, and Kaski (2012) Both the correlation parameter, $\rho_{s_{j}}$ and the specific component aim to mitigate negative transfer—learning behavior from secondary tasks that is not representative of the primary task. We make the prior assumption that $f_{p}$ and $\delta_{s_{j}}$ for $j=1,\dots,m$ follow Gaussian process distributions, each with a mean and kernel function which we will specify.
As was the case in subsections II.1 through II.3, the first step toward inference with the multitask method is to determine the joint distribution on observations and targets. Note that the specific component of each task is modeled as independent of the others. Then, the covariance between observations due to the specific component is a block matrix with zeros corresponding to the covariance of specific functions for different tasks. Call this block matrix $K^{\delta}$. Let $\bm{R}$ be a diagonal matrix with dimensions equal to the number of observations, and let each diagonal element be the correlation hyperparameter of the corresponding task. The covariance of the joint distribution is
@@
The superscript $p$ indicates that a matrix was created with the primary kernel, and the matrix subscripts indicate the features that are compared. Inference proceeds as described in Eq. (10).
II.5 Beyond Two Tasks
@@
$f_{s_{j}}(\bm{X})$ $=$ $\ \rho_{s_{j}}f_{p}(\bm{X})\ +\ \delta_{s_{j}}(\bm{X}).$
@@
Thus, compared to multifidelity fusion, the asymmetric multitasking approach more naturally accommodates a reasonably large number of regression tasks and relates each closely to the primary task of interest. Multitasking also holds a similar advantage over $\Delta$-learning. Before the notion of a difference can be imposed it is necessary to choose some order between the source data sets. Such an order may be arbitrary, for example, if our data sets are generated by multiple DFAs with similar structure. In Section III, we will describe different compositions of a training data set which the multitask framework allows us to test.
III Heterogeneous Training Data
@@
In our numerical examples, we train inference models using data for multiple molecular configurations computed with multiple quantum chemistry methods. The multitask method leaves many questions about the particulars of the training data set up to the researcher. For instance, available data must be apportioned to different tasks. One of our goals is to demonstrate that by training with data from multiple quantum chemistry methods, we can construct an inference model with prediction accuracy comparable to CCSD(T) with less computational expense than if we trained a single task model on CCSD(T) alone. Thus, our CCSD(T) data is the primary target of our regression problem and a natural choice to inform the primary task. We let each secondary task correspond to a different DFA. Examples with one up to four secondary tasks will be considered.
Another decision to be made is whether the training data sets for different tasks should include the same molecular configurations. In other words, we must determine whether or not to train with predictions for the same quantity by different methods. If our target for inference is the CCSD(T) prediction of energy for a system of water molecules, we may also choose to include DFT predictions of this energy in the training set for a secondary task. The relative benefits of constructing training sets with different degrees of intersection between tasks is under-explored in literature. Our work seeks to understand the impact of training set design on model performance.
We consider a range of possible training set compositions here, and Fig. 2 provides two example diagrams of our system for distinguishing between different constructions. We assign the molecular configurations to three sub data sets—Core (C), Additional (A), and Target (T). Dots in the figure indicate when configurations are available at a given level of theory. Black indicates availability to the training set, and green indicates membership of the testing set. The C set contains all configurations for which CCSD(T) training data is available, the A set contains configurations for which we have DFT data but no CCSD(T) data, and the T set contains configurations that we use to test our model’s ability to reproduce CCSD(T) level accuracy.

### v2

@@ II.3 $\Delta$-Learning
$\Delta$-learning performs Gaussian process regression to predict differences, $\Delta_{m}$, between observations obtained with two levels of theory for a given $\bm{X}_{m}$. In this setting, the observation data available to us allows two interpretations: either as a single-task approach with the observations $\{(\bm{Y}_{m2}-\bm{Y}_{m1})\}_{m=1}^{M}$, or as a two-task approach with data $\bigcup_{n=1}^{2}\{Y_{mn}\}_{m\in\mathcal{M}_{n}}$. Unless stated otherwise, we will assume the latter. If we retain our assumption that noise follows an independent, identically distributed Gaussian distribution, we have the model
[…]
$=$ $\ f_{\Delta}\big{(}\bm{X}_{m}\big{)}\ +\ \varepsilon_{m}$
$f_{\Delta}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{\Delta}(\bm{X}),\ k^{\Delta}(\bm{X},\bm{X}^{\prime})\big{)}.$
@@
where $n$ and $n^{\prime}$ indicate two prediction methods. An application of Eq. (4) yields a distribution on the regression function for $\Delta(n,n^{\prime})$. Suppose our ultimate interest is to predict $\bm{Y}$ at $\bm{X}_{*}$. Within the $\Delta$-learning framework, we can supply an observation from one method, say $\bm{Y}_{*n^{\prime}}$, to make a prediction for the value from the other method:
$\bm{Y}_{*n}\ \approx\ \Delta_{*}\ +\ \bm{Y}_{*n^{\prime}}.$
@@
Thus, $\Delta$-learning is suited to settings where calculations by method $n$ are both more accurate and more time consuming than calculations that use $n^{\prime}$.
@@
Consider a case where we want to build a model from an existing data set which consists of CCSD(T) predictions for configurations in $\mathcal{M}_{2}$ and DFT predictions for configurations in $\mathcal{M}_{1}$. In the notation introduced by Table 1, $\mathcal{M}_{2}$ is the core set, $C$. $\mathcal{M}_{1}$ may contain any combination of molecules from the $C$, $A$, and $T$ sets. As (5) indicates, the training data for $\Delta$-learning will include only molecular configurations in $\mathcal{M}_{1}\cap\mathcal{M}_{2}\subseteq C$. If $\mathcal{M}_{1}\cap\mathcal{M}_{2}=C$, then $\Delta$-learning will be able to leverage more data than the conventional GP regression model described by (3) trained on the CCSD(T) data. Unfortunately, $\Delta$-learning is not-applicable in cases where predictions by different methods correspond to different molecular configurations. When $\mathcal{M}_{1}\cap\mathcal{M}_{2}=\emptyset$, it is impossible to train a $\Delta$-learning model. To include a DFT prediction for a molecule $\bm{X}_{m}$ in the training set for a $\Delta$ model between CCSD(T) and DFT, we also require a CCSD(T) prediction for the same molecule.
Hierarchical machine learning (hML)Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) extends the $\Delta$-learning approach to $N>2$ by representing the difference between $\bm{Y}_{mN}$ and $\bm{Y}_{m1}$ as
$\Delta^{hML}_{m}(N,1)\ =\ \sum_{n=1}^{N-1}\Delta_{m}(n+1,n),$ (6)
where each $\Delta_{m}(n+1,n)$ is defined following (5) (with $f_{\Delta}$ potentially different for each $n$). As before, to ultimately predict, $\bm{Y}_{*N}$, we also require data for $\bm{Y}_{*1}$. Note that the summed difference models are independently trained, and their training sets need not be identical. Thus, an hML model designed to predict at the level of CCSD(T) would not necessarily require CCSD(T) data for every molecular configuration, $m$, in its training set. Rather, each training configuration must belong to some intersection of two adjacent levels of theory, i.e., to $\bigcup_{n=1}^{N-1}\left(\mathcal{M}_{n}\cap\mathcal{M}_{n+1}\right)$.
In the following subsections, we consider alternative approaches which model disparities in data sources and have the ability to use all available observations.
II.4 Multifidelity Method
@@
As a step toward considering statistical models which train on heterogeneous data, we describe the multifidelity data fusion model. Kennedy and O’Hagan (2000) The multifidelity approach defines two regression functions: $f_{p}$, representing the predictions from some high fidelity model, and $f_{s}$, representing a low fidelity model. Kennedy and O’Hagan (2000) We assume the relationships
$\bm{Y}_{mp}$ $=$ $\ f_{p}\big{(}\bm{X}_{m}\big{)}\ +\varepsilon_{mp},\qquad\forall\,m\in\mathcal{M}_{p}$ (7)
$=$ $\ \rho f_{s}\big{(}\bm{X}_{m}\big{)}\ +\ \delta_{ps}\big{(}\bm{X}_{m}\big{)}\ +\ \varepsilon_{mp},$
$\bm{Y}_{ms}$ $=$ $\ f_{s}\big{(}\bm{X}_{m}\big{)}\ +\ \varepsilon_{ms},\qquad\forall\,m\in\mathcal{M}_{s}.$ (8)
@@
The term $\delta_{ps}$ captures the difference between the high fidelity model and the low fidelity model, scaled by the hyperparameter $\rho$. This disparity is endowed with a Gaussian process prior, as is the low fidelity regression function, $f_{s}$:
$\delta_{ps}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{\delta}(\bm{X}),\ k^{\delta}(\bm{X},\bm{X}^{\prime})\big{)},$
$f_{s}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{s}((\bm{X})),\ k^{s}(\bm{X},\bm{X}^{\prime})\big{)}.$
The correlation, $\rho$, will be optimized before inference. As before, we assume that the additive noise terms, $\varepsilon_{mp}$ and $\varepsilon_{ms}$, are independent and drawn from identical, centered Gaussian distributions. Now, suppose we want to predict our QoI for the target molecules, $\{\bm{X}_{m}:m\in T\}$, at the primary fidelity. Let $\bm{f}_{*}$ be the vector of all values in $\{f_{p}(\bm{X}_{m}):m\in T\}$. We can determine that the joint distribution of our high and low fidelity observations and $\bm{f}_{*}$ is multivariate normal with mean and covariance
$\mu$ $=$ $\ \begin{bmatrix}\bm{\mu}^{s}\ +\ \bm{\mu}^{\delta}\\ \bm{\mu}^{s}\\ \bm{\mu}^{s}\ +\ \bm{\mu}^{\delta}\\ \end{bmatrix}$
@@
where $\sigma^{2}$ is the variance of the noise distribution. The superscript of each mean vector $\bm{\mu}$ and covariance matrix, K, corresponds to the GP prior used in its construction, and the subscripts indicate the pairs of features that are compared by the prior kernel function. Bayesian inference yields an analytical posterior distribution, analogous to Eq. (4).
II.5 Multitask Method
“Multitasking” refers to a category of methods which consider several regression tasks and assume some relationship between these tasks. As with $\Delta$-learning or multifidelity fusion, multitasking can allow us to incorporate a larger dataset into our inference problem without just pooling the data into one observation vector. For instance, we can define a regression problem trained on CCSD(T) data as well as one trained on each density-functional approximations (DFA) we consider. By modeling correlation between regression functions, we can use the DFA regression tasks to support prediction in the CCSD(T) task.
Symmetric variants of multitask modeling treat all tasks equally. For instance, we may assume that all the data are modeled by a single Gaussian process whose covariance kernel is the Kronecker product of one kernel which relates the tasks and another which models relationships between data from the same task.Bonilla, Chai, and Williams (2008); Khatamsaz, Vela, and Arróyave (2023) In many cases, however, the goal is to perform regression for a primary task, and all other tasks are considered secondary tasks which provide data that may be useful in learning the primary task. For our own setting, CCSD(T) is a natural choice of data for the primary regression task because we would like to make predictions which match this method in accuracy. Each set of DFT predictions made using a different DFA could inform a secondary regression task which supports the primary task. Thus, we consider an asymmetrical model Leen, Peltonen, and Kaski (2012) in which secondary tasks are related through the primary task.
Let $\mathcal{M}_{p}$ contain the molecular configurations available to train the primary task. Note that following the definitions provided in Table 1, $\mathcal{M}_{p}=C$. Likewise, we define $\mathcal{M}_{s_{n}}$ to be the set of training configurations for the $n^{th}$ secondary task. As before, we assume that each task has its own regression function
$\bm{Y}_{mp}\ =\ \ f_{p}(\bm{X}_{m})\ +\ \varepsilon_{mp},$ $\forall\,m\,\in\mathcal{M}_{p},$ (10)
$\bm{Y}_{ms_{n}}\ =\ f_{s_{n}}(\bm{X}_{m})\ +\ \varepsilon_{ms_{n}},$ $\forall\,m\,\in\mathcal{M}_{s_{n}},$
$\forall\,n=1,\dots,N-1.$
@@ II.5 Multitask Method
$f_{s_{n}}(\bm{X}_{m})$ $=\ \rho_{s_{n}}f_{p}(\bm{X}_{m})\ +\ \delta_{s_{n}}(\bm{X}_{m}).$ (12)
The task-specific component $\delta_{s_{n}}$ of the secondary function serves to “explain away” behavior that is not captured by the shared component, $f_{p}$een, Peltonen, and Kaski (2012) Both the correlation parameter $\rho_{s_{n}}$ and the task-specific component aim to mitigate negative transfer—learning behavior from secondary tasks that is not representative of the primary task. We make the prior assumption that $f_{p}$ and $\delta_{s_{n}}$ for $n=1,\dots,N-1$ follow Gaussian process distributions, each with a mean and kernel function which we will specify:
$f_{p}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{p}(\bm{X}),\ k^{p}(\bm{X},\bm{X}^{\prime})\big{)},$ (13)
$\delta_{s_{n}}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{s_{n}}(\bm{X}),\ k^{s_{n}}(\bm{X},\bm{X}^{\prime})\big{)}.$ (14)
As was the case in subsections II.1 through II.4, the first step toward inference of $\bm{f}_{*}$ with the multitask method is to determine the joint distribution of observations and targets. Since all secondary tasks are modeled following (12), we can easily represent this distribution for arbitrary $N$. Note that the task-specific component $\delta_{s_{n}}$ of each secondary task is modeled as independent of the others, $\delta_{s_{n^{\prime}}}$ for $n\neq n^{\prime}$. Then, the covariance between observations due to the tasks-specific components is a block diagonal matrix, which we denote as $K^{\delta}$. The kernel function $k^{s_{n}}$ is applied to construct the $s_{n}^{th}$ diagonal block. Since the off-diagonal blocks correspond to the covariances of task-specific specific functions for different tasks, they are all zero. We note that the number of rows and columns of this matrix both equal $\sum_{n=1}^{N-1}|\mathcal{M}_{s_{n}}|$. Let $\bm{R}$ be a diagonal matrix with dimensions equal to the number of secondary observations, and let each diagonal element be the correlation hyperparameter of the corresponding task. The mean and covariance of the multivariate normal joint distribution is
$\mu$ $=$ $\ \begin{bmatrix}\bm{\mu}^{p}\\ R\bm{\mu}^{p}+\bm{\mu}^{s}\\ \bm{\mu}^{p}\\ \end{bmatrix}$
@@
Superscripts indicate the GP prior structures used to create a vector or matrix, and subscripts indicate the features that are compared by a matrix. In both cases, $s$ indicates evaluations for all secondary tasks, $n=1,\dots,N-1$. Inference proceeds as described in Eq. (4).
II.6 Comparison of Methods for More than Two Fidelities
@@
$f_{s_{N}}(\bm{X})$ $=$ $\ \rho_{s_{N}}f_{p}(\bm{X})\ +\ \delta_{s_{N}}(\bm{X}).$
Thus, compared to multifidelity fusion, the asymmetric multitasking approach more naturally accommodates a reasonably large number of regression tasks and relates each closely to the primary task of interest.
@@
Multitasking also holds a similar advantage over the hML extension to $\Delta$-learning. The indices $n=1,\dots,N$ suggest an order of the training sets, and the difference models in (6) correspond to consecutive pairs of training sets. Such an order may be arbitrary, for example, if our data sets are generated by multiple DFAs with similar structure. In Section III, we will describe different compositions of a training data set which the multitask framework allows us to test.
III Mapping Heterogeneous Data to Tasks
$\mathcal{C}(t)\quad t\in\{p,s_{1},\dots,s_{N-1}\}$ intersection between core molecules and training molecules for task $t$
$\mathcal{A}(t)\quad t\in\{p,s_{1},\dots,s_{N-1}\}$ intersection between additional molecules and training molecules for task $t$
$\mathcal{T}(t)\quad t\in\{p,s_{1},\dots,s_{N-1}\}$ intersection between target molecules and training molecules for task $t$
Table 2: Definition of functions which return sets of molecular configurations corresponding to primary task, p, and secondary tasks, $s_{1},\dots,s_{N-1}$. Note that by definition, $\mathcal{C}(p)$ is the entire core set, $C$, and $\mathcal{A}(p)=\mathcal{T}(p)=\emptyset$.
@@ III Mapping Heterogeneous Data to Tasks
In our numerical examples, we train inference models using data for multiple molecular configurations computed with multiple quantum chemistry methods. The multitask method leaves many questions about the particulars of the training data set up to the researcher. In particular, it is necessary to define tasks and map molecules to the training set of each task.
III.1 Selecting Tasks
The primary task is chosen according to our prediction goal. We aim to demonstrate that by training with data from multiple quantum chemistry methods, we can construct an inference model which has prediction accuracy comparable to CCSD(T) for less computational expense than if we trained a single task model on CCSD(T) alone. Thus, we will test our surrogates using a set of CCSD(T) predictions for some target molecules belonging to a set we call $T$. Our primary task must be informed by CCSD(T) predictions for a different set of molecules which we refer to as our core training set, $C$. See Table 1 for an overview of our key definitions.
We consider two approaches to mapping between levels of theory and tasks. In the first (subsections IV.2.1 and IV.3), the training data of our primary task will consist of direct computations of our QoI at CCSD(T) level, and each secondary task will be trained by DFT predictions of the QoI obtained with different DFT functionals.
The second approach (Subsection IV.2.2) we call the “multitask $\Delta$” option. In that setting, the levels of theory include CCSD(T) and DFT computations performed with two different DFAs. The difference between CCSD(T) and the first DFA will be the primary task while the difference between the two DFAs informs the secondary task. Note that the $\Delta$-learning method could be interpreted as the single task counterpart to the multitask $\Delta$ option.
III.2 Assigning Molecular Configurations to Tasks
We are specifically interested in the flexibility which the multitask framework provides towards assigning molecular configurations to secondary training sets. Our pool of available molecules is divided into Core ($C$), Additional ($A$), and Target ($T$) subsets to systematically explore the impact of the molecular space covered by the secondary training data and its relationship to our primary training and testing data. The $A$ set contains all molecular configurations that are not part of the primary training or testing set. In our setting, these are configurations which are not our targets for inference and for which high level CCSD(T) predictions are unavailable. A major difference between the multitask framework and $\Delta$-learning is the former’s ability to leverage this $A$ set.
Fig. 1 shows the six main secondary training set structures considered in this work. We name the training set structures using combinations of the letters $C$, $A$, and $T$. Inclusion of a letter in a structure name indicates that at least one molecular configuration from the corresponding subset is included in the secondary training data set. For example, a CT training structure in our setting would contain CCSD(T) predictions for the C set of molecules, DFT predictions for some subset of the C set, and DFT predictions for some subset of the T set. The data requirement of implementing a multitask method with this CT training set structure is comparable to the data demands of a $\Delta$-learning model. However, a crucial distinction is that the multitask approach does not require a DFT prediction for every molecular configuration in the C and T sets. Furthermore, the multitask model can make use of an additional $A$ training set for which CCSD(T) predictions are unavailable. $\Delta$-learning models cannot be implemented for such a training set structure.
III.3 Multiple Secondary Tasks
The multitask framework can accommodate an arbitrary number of secondary tasks and allows the observational data used to train each task to correspond to non-identical sets of inputs. We will use the same definition for the $C$, $A$, and $T$ molecular sets regardless of the total number of tasks, $N$. In this context, when a training structure is labeled CAT, this name indicates that at least one configuration from each of these sets is present in the training set for at least one secondary task. To the best of our knowledge, most work which explores training set design for multifidelity, $\Delta$-learning, or hML models focuses on optimizing the size of the training set for each source. Questions of whether different secondary sources of information should explore the same or new areas of molecular space are underexplored in literature. For example, if our secondary training data includes DFT predictions at the PBE level for CO2 configurations and at the BLYP level for ONH4 configurations, we aim to determine how much our model will improve if we fill in the blanks (i.e., add PBE level predictions for ONH4 and BLYP level predictions for CO2 versus adding data on previously unseen molecules). Such questions are investigated in Subsection IV.3.